Finding the Equation of a Regression Line: Find the equation of the regression line for the data. Then construct a scatter plot of the data and draw the regression line. (Each pair of variables has a significant correlation.) Then use the regression equation to predict the value of y for each of the x-values, if meaningful. If the x-value is not meaningful to predict the value of y, explain why not. If convenient, use technology.

Square Footage and Home Sale Price: The square footages and sale prices (in thousands of dollars) of seven homes are shown in the table below.

  1. X = 1450 sq ft.   (b) X = 2720 sq ft. (c) X = 2175 sq ft. (d) X = 1890 sq ft.

X - Sq Ft.

Y - Sale Price

1924

174.9

1592

136.9

2413

275

2332

219.9

1552

120

1312

99.9

1287

145
Answer #1

Sol:

In excel select the data.

Go to

Insert>scatterchart

You will get

CHOLDATA - Excel Sao File Home Insert Page Layout Formulas Data Review View Help Tell me what you want to do 7 Get Add-ins My

Add axis titles,chart title

Click on one point

add trend line

nart 4 - x fc -SERIES(Sheet3!$B$1,Sheet3!$A$2:$A$8,Sheet3!$B$2:$B$8,1) A B C D E F G H I J K L M 9 300 Y - Sale X - Sq Ft. Pr

You get

y = 0.1234x - 51.412 R= 0.8581 Y - Sale Price 500 1000 1500 2000 2500 3000 X - Sq Ft.

Regression eq is

saleprice=0.1234*sqft-51.412

R sq=0.8581

=85.81% variation in sale price is explained by sq ft.

Good model.

when x= 1450 sq ft.

substitute in regression eq

y=0.1234*1450-51.412

y= 127.518

so sale price is 127.518

(b) X = 2720 sq ft

we cannot use this value as it is out of range for which we fit linear regression.

. (c) X = 2175 sq ft

y=0.1234*2175 -51.412

y= 216.983

when sq ft is 2175,sale price is 216.983

. (d) X = 1890 sq ft.

substitute in regression eq we get

y=0.1234*1890 -51.412

y= 181.814

when sq ft is 1890 ,sale price is 181.814